1. A mole of N2 has a pressure of 2.30 atm inside a 20.0 L vessel. What is the volume that the gas occupies if it is compressed to a pressure of 8.45 atm at constant temperature?

2.A sample of air in a container was analyzed to contain 0.390 mol N2, 0.102 mol O2 and 0.0085 mol other gases. What are the partial pressures of these gases in the container if the total pressure is 754 mm Hg?

3.What is the mass of ethanol that must be combusted to generate 20.4 L of CO2 at STP?

4.F2(g) (M=38.00 g/mol) effuses through a pinhole at a rate of 0.01287 mol/s. At what rate does N2O4(g) (M=92.02 g/mol) effuse at the same pinhole?


1
Expert's answer
2022-03-30T00:36:04-0400

1) "P1V1= P2V2"

2.3×20= 8.45×V2

V2= 46/8.45= 5.44L